# Heat dissipation in marginally stable linear time-delayed Langevin systems

### v3

@@ Discussion
In this work, we have studied the heat dissipation and its properties of two classes of criticality in marginally stable linear time-delayed Langevin systems. For diffusive criticality, despite its scaled diffusion, the probability distribution of heat dissipation becomes stationary and the average approaches a constant $aT\cdot\frac{1+a\tau}{1-a\tau}$ in the long-time limit. For oscillatory criticality, the distribution of heat dissipation continues to spread with time and the average asymptotically diverges as $\frac{4(b^{2}-a^{2})T}{1+b^{2}\tau^{2}-2a\tau}\cdot t+O(1)$ due to the diffusive amplitude of oscillation. These results show that the asymptotic behavior of the heat dissipation is determined by the characteristic roots of Eq. (5) with zero real parts for marginally stable linear time-delayed Langevin dynamics. Our results offer a reference point for further studies on stochastic thermodynamics of more general time-delayed systems, such as those under Pyragas control and near bifurcations.
One may face both significant challenges and fruitful insights when diving into the thermodynamics of the nonlinear time-delayed regime. A promising direction is to investigate stochastic resonance phenomena on nonlinear time-delayed systems. It is natural to expect that a nontrivial resonance may manifest in their heat dissipation profiles. Furthermore, nonlinearity may suppress the amplitude diffusion of bifurcations, leading to sustained finite-amplitude oscillations of time-delayed stochastic systems. Analyzing the thermodynamic properties of such oscillations would be another path for future research. Finally, the thermodynamic properties of collective behavior in time-delayed interacting stochastic systems remains an open area of research.
@@ Acknowledgements.
X. W. thanks Naruo Ohga, Sarah A. M. Loos, and Robin A. Kopp for enlightening discussions. X. W. acknowledges financial support from Department of Physics, Graduate School of Science, the University of Tokyo.
@@ Appendix A Numerical methods
We use the programming language Julia and the package DifferentialEquations.jl [47] and StochasticDelayDiffEq.jl [48] for numerical calculations and simulations (the Tsitouras 5/4 Runge-Kutta method with a free 4th order interpolant [49] is used for the numerical integration of $x_{0}(t)$ via Eq. (31); the Euler-Maruyama method [50] is used for trajectory simulations of stochastic dynamics). We draw the figures using the package Plots.jl [51] and GR Framework [52].
@@
We show how to derive Eq. (4) in detail based on the method in Ref. [39]. Consider the linear delayed differential equation (set $T=0$ in Eq. (1)):
@@
We expand the solution of Eq. (31) using an exponential series based on the roots of characteristic equation (Eq. (5)):
@@
where the coefficients are left to be determined according to the dynamics ($a,b$ and $\tau$) and the initial conditions ($\phi(t)$). To determine the coefficients, we first compute the Laplace transform of Eq. (31):
$sX(s)-x(0)=aX(s)+be^{-s\tau}X(s)+b\Phi(s),$ (34)
@@
$X(s)=\dfrac{x(0)+b\Phi(s)}{s-a-be^{-s\tau}}.$ (37)
@@
$C_{k}$ $=\underset{s=S_{k}}{\mathrm{Res}}\left(\dfrac{x(0)+b\Phi(s)}{s-a-be^{-s\tau}}\right)$
$=\dfrac{x(0)+b\Phi(S_{k})}{1+b\tau e^{-S_{k}\tau}}.$ (39)
@@
$x(t)=\sum_{k=-\infty}^{\infty}\dfrac{x(0)+b\Phi(S_{k})}{1+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (40)
@@
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{1}{1+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (41)

### v4

@@ Discussion
In this work, we have studied the heat dissipation and its properties of two classes of criticality in marginally stable linear time-delayed Langevin systems. For diffusive criticality, despite its scaled diffusion, the probability distribution of heat dissipation becomes stationary and the average approaches a constant $aT(\gamma+a\tau)/\gamma(\gamma-a\tau)$ in the long-time limit. For oscillatory criticality, the distribution of heat dissipation continues to spread with time and the average asymptotically diverges as $4(b^{2}-a^{2})T\cdot t/(\gamma^{2}+b^{2}\tau_{c}^{2}-2\gamma a\tau_{c})+O(1)$ due to the diffusive amplitude of oscillation. Notably, the constant asymptotic value of the average dissipation rate for diffusive criticality and the linearly divergence rate for oscillatory criticality are independent of the the history trajectory. We also analyze the asymptotic evolution of the heat dissipation rate as stable dynamics approaches these two critical boundaries, and find qualitative differences in both the convergence behavior and the associated spectral decomposition. This analysis demonstrates that non-stationary time-delayed dynamics can exhibit qualitatively different thermodynamic behaviors, depending crucially on the underlying dynamical details. Our results offer a reference point for further studies on stochastic thermodynamics of more general time-delayed systems, such as those under Pyragas control and near bifurcations.
One may face both significant challenges and fruitful insights when diving into the thermodynamics of the nonlinear time-delayed regime. A promising direction is to investigate stochastic resonance phenomena on nonlinear time-delayed systems. It is natural to expect that a nontrivial resonance may manifest in their heat dissipation profiles. Furthermore, nonlinearity may suppress the amplitude diffusion of Hopf bifurcations, leading to sustained finite-amplitude oscillations of time-delayed stochastic systems. Analyzing the thermodynamic properties of such oscillations would be another path for future research. The thermodynamic properties of collective behavior (e.g., synchronization or pattern formation) in many-body stochastic systems with time-delayed interaction remain an open area of research. Finally, a systematic comparison between the thermodynamics of time-delayed systems and other types of non-Markovian systems including active Brownian motion [58] and fractional Brownian motion [59] is also an interesting direction for future study.
@@ Acknowledgements.
X. W. thanks Sosuke Ito, Naruo Ohga, Sarah A. M. Loos, and Robin A. Kopp for enlightening discussions. X. W. acknowledges financial support from Department of Physics, Graduate School of Science, the University of Tokyo.
@@ Appendix A Numerical methods
We use the programming language Julia and the package DifferentialEquations.jl [60] and StochasticDelayDiffEq.jl [61] for numerical calculations and simulations (the Tsitouras 5/4 Runge-Kutta method with a free 4th order interpolant [62] is used for the numerical integration of $x_{0}(t)$ via Eq. (63); the Euler-Maruyama method [63] is used for trajectory simulations of stochastic dynamics). We draw the figures using the packages Plots.jl [64] and GR Framework [65].
@@
We show how to derive the exponential expansion of the fundamental solution [Eq. (4)] in detail using the method in Ref. [47]. Consider the linear delayed differential equation (set $T=0$ in Eq. (1)):
@@
We expand the solution of Eq. (63) using an exponential series based on the roots of characteristic equation [Eq. (5)]:
@@
where the coefficients are left to be determined according to the parameters $a,b$ and $\tau$ and the initial conditions $\phi(t)$. Note that we have assumed $\{S_{k}\}$ are all simple roots, which will be justified later. To determine the coefficients, we first compute the Laplace transform of Eq. (63):
${\color[rgb]{0,0,0}\gamma}[sX(s)-x(0)]=aX(s)+be^{-s\tau}X(s)+b\Phi(s),$ (66)
@@
$X(s)=\dfrac{{\color[rgb]{0,0,0}\gamma}x(0)+b\Phi(s)}{{\color[rgb]{0,0,0}\gamma}s-a-be^{-s\tau}}.$ (69)
@@
$C_{k}$ $=\underset{s=S_{k}}{\mathrm{Res}}\left(\dfrac{{\color[rgb]{0,0,0}\gamma}x(0)+b\Phi(s)}{{\color[rgb]{0,0,0}\gamma}s-a-be^{-s\tau}}\right)$
$=\dfrac{{\color[rgb]{0,0,0}\gamma}x(0)+b\Phi(S_{k})}{{\color[rgb]{0,0,0}\gamma}+b\tau e^{-S_{k}\tau}}.$ (71)
@@
$x(t)=\sum_{k=-\infty}^{\infty}\dfrac{{\color[rgb]{0,0,0}\gamma}x(0)+b\Phi(S_{k})}{{\color[rgb]{0,0,0}\gamma}+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (72)
@@
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{{\color[rgb]{0,0,0}\gamma}}{{\color[rgb]{0,0,0}\gamma}+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (73)
We now show that for both diffusive and oscillatory criticality, the denominator of the coefficient $\gamma/(\gamma+b\tau e^{-S_{k}\tau})$ is nonzero, i.e. all roots of characteristic equation Eq. (5) are simple. If a root were not simple, it must satisfy both:
$\gamma S-a-be^{-S\tau}=0,$ (74)
$\gamma+b\tau e^{-S\tau}=0,$ (75)
which yields
$b=-\frac{\gamma}{\tau}e^{\frac{a\tau}{\gamma}-1}.$ (76)
For the diffusive criticality, $a+b=0$ as in Eq. (8); thus $a=(\gamma/\tau)e^{a\tau/\gamma-1}$. Setting $x=a\tau/\gamma$, we have
$x=e^{x-1},$ (77)
which has a unique solution $x=1$, implying $a\tau=\gamma$. This contradicts the assumption $a\tau<\gamma$ in Eq. (8). For oscillatory criticality ($\tau=\tau_{c}$), from Eqs. (15) and setting $\omega_{c}\tau=\eta$, we have
$a\tau_{c}=\gamma\eta\cot\eta,\ \ b\tau_{c}=-\frac{\gamma\eta}{\sin\eta}.$ (78)
Combining these expressions with Eq. (76), we obtain
$e^{\eta\cot\eta-1}=\frac{\eta}{\sin\eta}.$ (79)
Since $\eta=\arccos(-a/b)\in(0,\pi)$ in Eq. (13), $\sin\eta<\eta$ while $\eta\cot\eta<1$; therefore oscillatory criticality does not admit roots that are not simple.
Appendix C Proof of the boundedness of $C_{0}$ and $C_{1}$
We explicitly prove the boundedness of the constants $C_{0}$ and $C_{1}$ in Eqs. (12) and (18). We consider the overdamped time-delayed marginally stable dynamics, where the fundamental solution $x_{0}(t)$ satisfies
$x_{0}(t)=\sum_{k=-\infty}^{\infty}\dfrac{\gamma}{\gamma+b\tau e^{-S_{k}\tau}}e^{S_{k}t}\ \ (t>0),$ (80)
and $\{S_{k}\}\ (k\in\mathcal{Z})$ are all roots of the characteristic equation [47]
$\gamma S_{k}-a-be^{-S_{k}\tau}=0.$ (81)
Subtracting from $x_{0}(t)$ the terms corresponding to the characteristic roots with zero real parts, we define :
$x_{0,s}(t)$ $=x_{0}(t)-\sum_{\mathrm{Re}[S_{k}]=0}\dfrac{\gamma}{\gamma+b\tau e^{-S_{k}\tau}}e^{S_{k}t}$
$=\sum_{\mathrm{Re}[S_{k}]<0}\dfrac{\gamma}{\gamma+b\tau e^{-S_{k}\tau}}e^{S_{k}t}.$ (82)
In the following, we show the absolute convergence of $x_{0,s}(t)$ for $t>0$. Convergence at $t=0$ is ensured by the initial condition. Using Eqs. (C) and (81), we obtain
$x_{0,s}(t)=\sum_{\mathrm{Re}[S_{k}]<0}\dfrac{\gamma}{\gamma-a\tau+\gamma S_{k}}e^{S_{k}t}.$ (83)
From the characteristic equation (81), we have
$e^{S_{k}\tau}=\frac{b}{\gamma S_{k}-a}=\frac{b}{\gamma S_{k}}\cdot\frac{1}{1-a/\gamma S_{k}}.$ (84)
Taking the logarithm and expanding for large $|S_{k}|{\color[rgb]{0,0,0}\gg|a|/\gamma}$, we obtain
$S_{k}\tau=2\pi ik+\ln b-\ln\gamma-\ln S_{k}+\frac{a}{\gamma S_{k}}+O(|S_{k}|^{-2}),\ \ k\in\mathbb{Z}.$ (85)
Let $S_{k}=\sigma_{k}+i\lambda_{k}$. From Eq. (85), the real and imaginary parts are
$\sigma_{k}\tau=\ln|b|-\ln\gamma-\ln|S_{k}|+\operatorname{Re}\left(\frac{a}{\gamma S_{k}}\right)+O(|S_{k}|^{-2}),$ (86)
$\lambda_{k}\tau=2\pi k+\arg b-\arg S_{k}+\operatorname{Im}\left(\frac{a}{\gamma S_{k}}\right)+O(|S_{k}|^{-2}),$ (87)
respectively (for a more precise asymptotic analysis, see Ref. [66]). Here, we take $\arg b=0$ for positive $b$, $\arg b=\pi$ for negative $b$, and $\arg S\in(-\pi,\pi]$. For large $|k|$, Eqs. (86) and (87) give
$|\lambda_{k}|$ $\sim\frac{2\pi|k|}{\tau}+{\color[rgb]{0,0,0}M_{0}}+O(|S_{k}|^{-1}),$ (88)
$\sigma_{k}$ $\sim-\frac{1}{\tau}\ln|k|+{\color[rgb]{0,0,0}K_{0}}+O(|S_{k}|^{-1}),$ (89)
${\color[rgb]{0,0,0}|S_{k}|}$ ${\color[rgb]{0,0,0}=|\lambda_{k}|\sqrt{1+\sigma_{k}^{2}/\lambda_{k}^{2}}\sim|\lambda_{k}|\cdot\bigg[1+O\bigg((\ln|k|/k)^{2}\bigg)\bigg],}$ (90)
where ${\color[rgb]{0,0,0}M_{0}}\equiv\arg b-\arg S\in(-2\pi,2\pi]$ and $K_{0}\equiv{\color[rgb]{0,0,0}\ln(|b|\tau/2\pi\gamma)}$. Therefore, there exists a sufficient large $N_{0}>0$, such that when $|k|>N_{0}$, the relations at below hold:
$|\lambda_{k}|$ $\geq\frac{2\pi|k|}{\tau}-|M_{0}|-1,$ (91)
$\sigma_{k}$ $\leq-\frac{1}{\tau}\ln|k|+|K_{0}|+1,$ (92)
$|S_{k}|$ $\geq|\lambda_{k}|/2.$ (93)
Taking $N_{1}\equiv[(|M_{0}|+1)\tau/2\pi]+1$ and noting that $a\tau<\gamma$, we have
$\Biggl|\dfrac{\gamma}{\gamma-a\tau+\gamma S_{k+N_{1}}}\Biggr|<\frac{\tau}{\pi}\cdot\frac{1}{|k|}.$ (94)
We split $x_{0,s}(t)$ for $t>0$ into two parts:
$x_{0,s}(t)=\sum_{|k|\leq{\color[rgb]{0,0,0}N_{0}}+N_{1},\ \mathrm{Re}[S_{k}]<0}\dfrac{\gamma}{\gamma-a\tau+\gamma S_{k}}e^{S_{k}t}$
$+\sum_{|k|>{\color[rgb]{0,0,0}N_{0}}+N_{1},\ \mathrm{Re}[S_{k}]<0}\dfrac{\gamma}{\gamma-a\tau+\gamma S_{k}}e^{S_{k}t}$ (95)
For a specified $N_{0}$, the first sum contains finite terms; hence there exists a constant $M_{1}$ such that
$\Biggl|\sum_{|k|\leq{\color[rgb]{0,0,0}N_{0}}+N_{1},\ \mathrm{Re}[S_{k}]<0}\dfrac{\gamma}{\gamma-a\tau+\gamma S_{k}}e^{S_{k}t}\Biggr|\leq|M_{1}|e^{-\beta t},$ (96)
where $\beta=\underset{\mathrm{Re}[S_{k}]<0}{\mathrm{Inf}}\Big|\mathrm{Re}[S_{k}]\Big|$. For the second sum, using Eqs. (92) and  (94):
$\Bigg|\sum_{|k|>{\color[rgb]{0,0,0}N_{0}}+N_{1},\ \mathrm{Re}[S_{k}]<0}\dfrac{\gamma}{\gamma-a\tau+\gamma S_{k}}e^{S_{k}t}\Bigg|$
$\leq\frac{\tau}{\pi}e^{({\color[rgb]{0,0,0}|K_{0}|}+1)t}\sum_{|k|\geq{\color[rgb]{0,0,0}N_{0}}}|k|^{-(1+t/\tau)}$
$\leq\frac{2\tau}{\pi}e^{({\color[rgb]{0,0,0}|K_{0}|}+1)t}\int_{{\color[rgb]{0,0,0}N_{0}}}^{\infty}x^{-(1+t/\tau)}dx$
$=\frac{2\tau^{2}}{\pi t}e^{({\color[rgb]{0,0,0}|K_{0}|}+1-\ln{\color[rgb]{0,0,0}N_{0}}/\tau)t}.$ (97)
Here we have used the condition $a\tau<\gamma$. Choosing $N_{0}$ large enough so that $|K_{0}|+1-\ln N_{0}/\tau<-\beta$, then Eqs. (96) and (C) imply the existence of a constant $M>0$ such that
$|x_{0,s}(t)|\leq M(1+t^{-1})e^{-\beta t},\ \ \ \ \ \text{for}\ t>0.$ (98)
This verifies the boundedness of $x_{0,s}(t)$ for $t>0$ and shows for large times $t$,
$x_{0,s}(t)=O(e^{-\beta t}).$ (99)
By a similar procedure, we can also show that for the fundamental solution $x_{0}(t)$ of the stable dynamics [Eq. (1)] under the condition Eq. (30) or (31), there exists a constant $M_{2}>0$ such that
$|x_{0}(t)|\leq M_{2}(1+t^{-1})e^{-\beta t},\ \ \ \ \ \text{for}\ t>0.$ (100)
For a sufficiently large time,
$x_{0}(t)=O(e^{-\beta t}).$ (101)
Now consider the constant $C_{0}$:
$C_{0}$ $=\frac{2T}{\gamma}\int_{0}^{\infty}\left[x_{0}^{2}(s)-\left(\dfrac{\gamma}{\gamma-a\tau}\right)^{2}\right]ds$
$=\frac{2T}{\gamma}\int_{0}^{\infty}x_{0,s}(s)\Bigg[x_{0,s}(s)+\dfrac{2\gamma}{\gamma-a\tau}\Bigg]ds.$ (102)
The convergence speed of $C_{0}$ is determined by that of $x_{0,s}(t)$ and is also comparable to $O(e^{-\beta t})$. Equation (98) ensures the boundedness on any interval $(T_{0},\infty]$ with $T_{0}>0$ . Combined with the continuity and hence boundedness of $x_{0}(t)$ on $(0,T_{0}]$, the integral in Eq. (102) is indeed bounded. A similar procedure applies to $C_{1}$ in Eq. (18).